What is the effective weight of a person of mass 75 kg carried vertically up in a rocket with an acceleration of 1.5g?


1
Expert's answer
2021-01-22T17:00:51-0500
"N-mg=ma,""N=m(g+a)=75\ kg\cdot(9.8\ \dfrac{m}{s^2}+1.5\cdot9.8\ \dfrac{m}{s^2})=1837.5\ N."